Jorge wishes to make 150 pounds of coffee blend that can be sold for $8 per pound. The blend will be a mixture of coffee worth $6 per pound and coffee worth $12 per pound. How many pounds of each kind of cofee whould be used in the mixture?
Answer by checkley77(12844)

12x+6(150-x)=8*150
12x+900-06x=1,200
6x=1,200-900
6x=300
x=300/6
x=50 lbs. of $12 coffee is used.
150-50=100 lbs. of $6 coffee is used.
Proof:
12*50+6*100=1,200
600+600=1,200
1,200=1,200